### v1

@@ 2.2 Angle-averaged cross sections
In order to estimate the impact of DM self-scattering on various observables of interest, it is useful to define integrated, angle-averaged cross sections with specific weighting factors. These may also be used to map results obtained for a given angle-dependence (e.g. isotropic) to other models, if the effect of interest is mainly sensitive to a particular averaged cross section. Here we briefly review the angle-averages that are often considered in the context of SIDM. We use them below to (i) characterise the relative strength of small- and large-angle scatterings in a given model, and (ii) investigate in how far our results taking the precise angle-dependence into account can indeed be mapped among models featuring distinct differential cross sections, but agreeing in the value of a certain angle-averaged cross section.
@@
The viscosity cross sections has been argued to be a good proxy for the impact of self-scattering on the distribution of DM in isolated halos, including core formation and collapse (e.g. Colquhoun et al., 2021; Yang & Yu, 2022; Sabarish et al., 2024). The question whether an average cross section is a good description in mergers has also been investigated (e.g. Kahlhoefer et al., 2014; Robertson et al., 2017a; Fischer et al., 2021a, b). Below we scrutinise these findings for the evolution of DM and galaxy positions as well as their offset in cluster mergers, by comparing to a treatment capturing the exact angle-dependence. The total cross section is mainly relevant for estimating the numerical effort (within the rSIDM scheme, see below) since it characterises the total scattering rate, irrespective of whether a significant amount of momentum is transferred between the scattering partners.
@@
2.3 Existing approaches for angle-dependent SIDM
Here we briefly review the physical basis of the approaches that have been used to incorporate angle-dependence into SIDM simulations, while technical details on the numerical implementation are provided in Sect. 4.
@@ 2.3 Existing approaches for angle-dependent SIDM
The first method treats angle-dependence explicitly for each scattering event, by sampling the deflection angle from a random distribution that follows from the differential cross section, see e.g. Robertson et al. (2017a); Banerjee et al. (2020). This strategy is in practice numerically feasible for models for which an ${\cal O}(1)$ amount of energy and momentum is transferred in each scattering, being thus limited to large-angle scatterings, see Robertson et al. (2017a). For these types of models the total and (modified) transfer cross sections are of comparable size, $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\sim{\cal O}(1)$ which implies that even relatively ‘rare’ scattering events can have a sizeable impact. We thus refer to this case as rare self-interacting dark matter (rSIDM).
@@
The second method employs an effective treatment that is applicable if DM scatters very frequently, but the momentum transfer in each scattering is small, dubbed frequent self-interacting dark matter (fSIDM) following Fischer et al. (2021a). By itself, this method becomes exact for models approaching the formal limit $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\to\infty$, that is very strongly forward-dominated scattering only.
@@
In this work we combine both methods, using them within the range of angles for which they work best. Before discussing this hybrid approach, we briefly review the physical basis of the fSIDM scheme, see Kahlhoefer et al. (2014) and Kummer et al. (2018) for more details. In analogy to the seminal description of Brownian motion by Einstein (1905) and von Smoluchowski (1906), the collective effect of frequent small-angle scattering can be described by a drag force that acts along the direction of motion (more precisely the relative velocity vector) as well as a stochastic force leading to diffusive motion. The drag force can be related to the (modified) transfer cross section, $F_{\rm drag}\propto\sigma_{\mathrm{\widetilde{T}}}$ (see Sect. 4 for details), while the stochastic force can be included as a kick in a random direction transverse to the relative velocity vector, with magnitude consistent with momentum and energy conservation.
@@
This set-up has been implemented by Fischer et al. (2021a) and verified with multiple test problems including one for which the solution is known from Molière (1948), being essentially a Gaussian spreading of an initially collimated beam traversing a medium of target particles. In this work we generalise this test problem to models with arbitrary differential cross section in order to validate the hybrid scheme based on combining the rSIDM and fSIDM methods, to which we turn next.
3 Hybrid scheme for angle-dependence
@@
In this section we first present an algorithm allowing us to simulate DM self-scattering with arbitrary differential cross sections efficiently, dubbed hybrid self-interacting dark matter or hSIDM. Next, we review a test set-up featuring an exact analytical solution for any given differential cross section, and then use it to assess under which conditions and for which choices of various technical parameters the hSIDM approach is expected to be valid.
@@
3.1 Hybrid approach to angle-dependent SIDM
@@
Here we introduce a new hybrid approach for taking the dependence of the DM self-interaction cross section on the deflection angle into account, that we refer to as hSIDM as stated above. The central idea is to combine the methods used previously to describe models with either only large (rSIDM) or only forward-dominated (fSIDM) scatterings, allowing for both regimes to be described efficiently within a single set-up. Since typical (e.g. light mediator) models can feature a very strong enhancement of the total scattering rate at small angles, see e.g. Fig. 1, but also predict non-negligible large-angle scattering, it is necessary to capture both contributions in order to obtain accurate predictions for a given set of model parameters.
@@
Since the critical angle $\theta_{\mathrm{c}}$ is a purely technical parameter, all physical quantities should be independent of its choice, and we check this explicitly for all our results below. Theoretically, we expect this is the case provided $\theta_{\mathrm{c}}$ is chosen small enough such that the small-angle effective description underlying the fSIDM approach can be applied. As we see later on, the appropriate size of $\theta_{\mathrm{c}}$ depends on characteristics of the set-up, such as the DM density and the considered time interval. We quantify the validity range of the hSIDM scheme in Sect. 3.2, employing a test set-up that can be described analytically.
Figure 2: Ratio of the contribution to the total cross section from the large-angle regime (dashed lines) and of the modified transfer cross section from the small-angle regime, versus the critical angle $\theta_{\mathrm{c}}$ used to separate small- and large-angle regions within the hSIDM scheme, for the ‘Møller’ cross section Eq. (1) and $r=10^{2},10^{3},10^{4}$, respectively. Solid lines characterise the expected physical relevance of small-angle scattering (in relative units). For moderate values of $\theta_{\mathrm{c}}\lesssim 0.5$, dashed lines conservatively estimate the expected ratio of the numerical effort of hSIDM compared to a naive implementation of angle-dependent scattering.
@@
The numerical costs depend on the size of the numerical time-step. Within the fSIDM scheme, the time-step is proportional to the drag force, $F_{\rm drag}$, and hence also proportional to the (modified) transfer cross section, $\sigma_{\mathrm{\widetilde{T}}}$. In contrast, the rSIDM scheme, for which individual collisions are modelled numerically as scattering of $N$-body particles (see Sect. 4), has a time-step proportional to the total scattering rate, $R$, which is proportional to the total cross section of large-angle scattering, i.e. within the hSIDM scheme it is
@@
We note that for light mediator models (i.e. ‘Møller scattering’), the strong enhancement of the forward-scattering rate implies that for typical values of $\theta_{\mathrm{c}}\sim 0.1-0.3$ one has $\sigma_{{\rm tot},>}(\theta_{\mathrm{c}})\ll\sigma_{\rm tot}$. The expected speed-up, $S$, within hSIDM as compared to a naive direct treatment of the angle-dependence can thus be estimated as
@@
In summary, the main advantage of the hSIDM regime is that it allows us to treat very frequent small-angle scatterings much more efficiently as compared to a naive sampling of deflection angles, while in addition explicitly including the angle-dependence of the physically roughly equally important contribution from large-angle scattering. With the hSIDM scheme simulation of fairly anisotropic cross sections with significant contributions from both small and large angles that have not been feasible before, become possible thanks to the efficient treatment of small-angle scattering.
@@ 3.2 Applicability of effective description for small-angle scattering: deflection test set-up
In order to gain insight on the expected range of validity of the hSIDM scheme, and especially on the choice of the critical angle $\theta_{\mathrm{c}}$, we investigate the applicability of the effective treatment of small-angle scattering. For that purpose, it is instructive to consider a simple test set-up that admits an analytical treatment.
@@ 3.2.1 Deflection test set-up
The approximation given by Eq. (24) is the essential assumption underlying the effective treatment of small-angle scattering entering the fSIDM approach. This can be seen more directly when assuming in addition that $\tau_{\mathrm{T}}$ is small enough such that sufficiently many $\ell$ contribute in the summation in Eq. (25). In this case the summation over $\ell$ can be replaced by a continuous integration using the Euler formula, see e.g. Bethe (1953), and Eq. (25) yields the Molière approximation (compare with Molière, 1948)
@@ 3.2.2 Validity of the fSIDM approach
Importantly, the analytically solvable deflection set-up allow us to derive a general criterion for the validity of the effective forward-dominated approximation. In this subsection, we start by discussing the validity in the case of allowing for the entire range of possible scattering angles, i.e. the fSIDM approximation. Subsequently, we generalise the discussion to the hSIDM case, for which only angles below a critical angle are included in the effective small-angle approximation.
@@
where we expressed the number density in terms of the energy density using $\rho=m_{\chi}n$. For the effective small-angle approach to be useful, $t_{\rm min}$ should be much smaller than the time-scale on which the distribution reaches its late-time isotropic limit, which is of order of the typical dynamical evolution time-scale $t_{\rm dyn}\equiv 1/(\rho v\frac{\sigma_{\mathrm{T}}}{m_{\chi}})$, where $v$ is the beam velocity. Thus, we arrive at the condition for the effective small-angle treatment to be applicable and useful, given by
@@
We note that this is the condition that is relevant when using the effective fSIDM approach for all possible deflection angles, being in practice limited to models that exclusively feature forward-dominated scattering. For realistic models, this is typically not the case. As an example, Møller scattering with $r=1,10,100,10^{3},10^{4}$ yields $\sigma_{\mathrm{T}^{2}}/\sigma_{\mathrm{T}}=0.32,0.23,0.14,8.5\times 10^{-2},5.9\times 10^{-2}$ respectively. Analogously Rutherford scattering yields $1.2,0.82,0.51,0.33,0.24$. This motivates using the hybrid approach instead. We discuss its validity next.
@@ 3.2.3 Validity of the hSIDM approach
where $\tau_{X,<}(t,\theta_{\mathrm{c}})\equiv nvt\sigma_{X,<}(\theta_{\mathrm{c}})$ for $X=\mathrm{T}$ (transfer cross section) and $X=T^{2}$ (transfer-squared cross section, defined analogously to Eq. (29) but integrated only over the angle-interval $I_{<}$). As above, this implies a minimal time-scale after which the hSIDM approach is valid, given by
@@
The conditions from above can be easily generalised to the case of identical particles with equal mass by using the appropriate definition of the cross sections $\sigma_{X,<}(\theta_{\mathrm{c}})$ for that case, (a) integrated over the interval corresponding $I_{<}=[0,\theta_{\mathrm{c}}]\cup[\uppi-\theta_{\mathrm{c}},\uppi]$ instead of $I_{<}=[0,\theta_{\mathrm{c}}]$, and (b) replacing the transfer ($X=\rm T$) by the modified transfer ($X=\rm{\widetilde{T}}$) cross section, see Sec. 2.2, and analogously for $X=\mathrm{T}^{2}$.
As an illustrative example, for Møller (Rutherford) scattering with $r=10^{3}$, one obtains $\sigma_{\mathrm{T}^{2},<}(\theta_{\mathrm{c}})/\sigma_{\mathrm{T},<}(\theta_{\mathrm{c}})=1.3\times 10^{-3}\,(2.6\times 10^{-3})$ for $\theta_{\mathrm{c}}=0.1$. We provide a detailed comparison of hSIDM simulations to a full, explicit treatment of the angle-dependence for various models in Sect. 5, where we also cross-check the theoretically expected validity conditions derived here.

### v2

@@ 2.2 Angle-averaged cross sections
In order to estimate the impact of DM self-scattering on various observables of interest, it is useful to define integrated, angle-averaged cross sections with specific weighting factors. These may also be used to map results obtained for a given angle dependence (e.g. isotropic) to other models, if the effect of interest is mainly sensitive to a particular averaged cross section. Here we briefly review the angle-averages that are often considered in the context of SIDM. We use them below to (i) characterise the relative strength of small- and large-angle scatterings in a given model, and (ii) investigate in how far our results taking the precise angle dependence into account can indeed be mapped among models featuring distinct differential cross sections, but agreeing in the value of a certain angle-averaged cross section.
@@
The viscosity cross sections has been argued to be a good proxy for the impact of self-scattering on the distribution of DM in isolated halos, including core formation and collapse (e.g. Colquhoun et al., 2021; Yang & Yu, 2022; Sabarish et al., 2024). The question whether an average cross section is a good description in mergers has also been investigated (e.g. Kahlhoefer et al., 2014; Robertson et al., 2017a; Fischer et al., 2021a, b). Below we scrutinise these findings for the evolution of DM and galaxy positions as well as their offset in cluster mergers, by comparing to a treatment capturing the exact angle dependence. The total cross section is mainly relevant for estimating the numerical effort (within the rSIDM scheme, see below) since it characterises the total scattering rate, irrespective of whether a significant amount of momentum is transferred between the scattering partners.
@@
2.3 Existing approaches for angle-dependent self-interacting dark matter
Here we briefly review the physical basis of the approaches that have been used to incorporate angle dependence into SIDM simulations. Technical details on the numerical implementation are provided in Sect. 4.
@@ 2.3 Existing approaches for angle-dependent self-interacting dark matter
The first method treats angle dependence explicitly for each scattering event, by sampling the deflection angle from a random distribution that follows from the differential cross section (see e.g. Robertson et al., 2017a; Banerjee et al., 2020). This strategy is in practice numerically feasible for models for which an ${\cal O}(1)$ amount of energy and momentum is transferred in each scattering, thus being limited to large-angle scatterings, see Robertson et al. (2017a). For these types of models the total and (modified) transfer cross sections are of comparable size, $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\sim{\cal O}(1)$ , which implies that even relatively ‘rare’ scattering events can have a sizeable impact. We thus refer to this case as rare self-interacting dark matter (rSIDM).
@@
The second method employs an effective treatment that is applicable if DM scatters very frequently, but the momentum transfer in each scattering is small, dubbed frequent self-interacting dark matter (fSIDM) following Fischer et al. (2021a). By itself, this method becomes exact for models approaching the formal limit $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\to\infty$, that is, very strongly forward-dominated scattering only.
@@
In this work we combine both methods, using them within the range of angles for which they work best. Before discussing this hybrid approach, we briefly review the physical basis of the fSIDM scheme (see Kahlhoefer et al., 2014; Kummer et al., 2018, for more details). In analogy to the seminal description of Brownian motion by Einstein (1905) and von Smoluchowski (1906), the collective effect of frequent small-angle scattering can be described by a drag force that acts along the direction of motion (more precisely the relative velocity vector) as well as a stochastic force leading to diffusive motion. The drag force can be related to the (modified) transfer cross section, $F_{\rm drag}\propto\sigma_{\mathrm{\widetilde{T}}}$ (see Sect. 4 for details), while the stochastic force can be included as a kick in a random direction transverse to the relative velocity vector, with magnitude consistent with momentum and energy conservation.
@@
This set-up has been implemented by Fischer et al. (2021a) and verified with multiple test problems including one for which the solution is known from Molière (1948), being essentially a Gaussian spreading of an initially collimated beam traversing a medium of target particles. In this work we generalise this test problem to models with arbitrary differential cross section in order to validate the hybrid scheme based on combining the rSIDM and fSIDM methods, which we turn to next.
3 Hybrid scheme for angle dependence
@@
In this section, we first present an algorithm that allows us to simulate DM self-scattering with arbitrary differential cross sections efficiently, dubbed hybrid self-interacting dark matter (hSIDM). Next, we review a test set-up featuring an exact analytical solution for any given differential cross section and then use it to assess under which conditions and for which choices of various technical parameters the hSIDM approach is expected to be valid.
@@
3.1 Hybrid approach to angle-dependent self-interacting dark matter
@@
Here we introduce a new hybrid approach for taking the dependence of the DM self-interaction cross section on the deflection angle into account, that we refer to as hSIDM as stated above. The central idea is to combine the methods used previously to describe models with either only large (rSIDM) or only forward-dominated (fSIDM) scatterings, allowing for both regimes to be described efficiently within a single set-up. Since typical (e.g. light mediator) models can feature a very strong enhancement of the total scattering rate at small angles, (see e.g. Fig. 1) but also predict non-negligible large-angle scattering, it is necessary to capture both contributions in order to obtain accurate predictions for a given set of model parameters.
@@
Since the critical angle $\theta_{\mathrm{c}}$ is a purely technical parameter, all physical quantities should be independent of its choice, and we check this explicitly for all our results below. Theoretically, we expect this is the case provided $\theta_{\mathrm{c}}$ is chosen small enough such that the small-angle effective description underlying the fSIDM approach can be applied. As we see later on, the appropriate size of $\theta_{\mathrm{c}}$ depends on characteristics of the set-up, such as the characteristic dynamical time and the considered time interval, in addition to the angle dependence of the cross section itself. We quantify the validity range for the choice of $\theta_{\mathrm{c}}$ within the hSIDM scheme in Sect. 3.2, employing a test set-up that can be described analytically. In addition, we point to Sect. 5.3 for a discussion of requirements constraining the choice of $\theta_{\mathrm{c}}$ within typical simulation set-ups, and to Sect. 5.4 for a demonstration of the performance gain of the hybrid scheme compared to a direct implementation of angular dependence.
Figure 2: Ratio of the contribution to the total cross section from the large-angle regime (dashed lines) and of the modified transfer cross section from the small-angle regime versus the critical angle $\theta_{\mathrm{c}}$ used to separate small- and large-angle regions within the hSIDM scheme for the ‘Møller’ cross section Eq. (1) and $r=10^{2},10^{3},10^{4}$, respectively. Solid lines characterise the expected physical relevance of small-angle scattering (in relative units). For moderate values of $\theta_{\mathrm{c}}\lesssim 0.5$, dashed lines conservatively estimate the expected ratio of the numerical effort of hSIDM compared to a naive implementation of angle-dependent scattering.
@@
The numerical costs depend on the size of the numerical time step. Within the fSIDM scheme, the time step is proportional to the drag force, $F_{\rm drag}$, and hence also proportional to the (modified) transfer cross section, $\sigma_{\mathrm{\widetilde{T}}}$. In contrast, the rSIDM scheme, for which individual collisions are modelled numerically as scattering of $N$-body particles (see Sect. 4), has a time step proportional to the total scattering rate, $R$, which is proportional to the total cross section of large-angle scattering, i.e. within the hSIDM scheme it is
@@
We note that for light mediator models (i.e. ‘Møller scattering’), the strong enhancement of the forward-scattering rate implies that for typical values of $\theta_{\mathrm{c}}\sim 0.1-0.3$ one has $\sigma_{{\rm tot},>}(\theta_{\mathrm{c}})\ll\sigma_{\rm tot}$. The expected speed-up, $S$, within hSIDM as compared to a naive direct treatment of the angle dependence can thus be estimated as
@@
In summary, the main advantage of the hSIDM regime is that it allows us to treat very frequent small-angle scatterings much more efficiently as compared to a naive sampling of deflection angles, while in addition explicitly including the angle dependence of the physically roughly equally important contribution from large-angle scattering. With the hSIDM scheme simulation of fairly anisotropic cross sections with significant contributions from both small and large angles that have not been feasible before, become possible thanks to the efficient treatment of small-angle scattering.
@@ 3.2 Applicability of effective description for small-angle scattering: deflection test set-up
In order to gain insight on the expected range of validity of the hSIDM scheme, and especially on the choice of the critical angle $\theta_{\mathrm{c}}$, we investigate the applicability of the effective treatment of small-angle scattering. For that purpose, it is instructive to consider a simple test set-up that admits an analytical treatment. The purpose of considering this set-up is $(i)$ to derive analytical conditions on the angle dependence of the cross section, as well as the typical local dark matter density and velocity and dynamical timescale for which the effective treatment of small-angle scattering is valid (see Sect. 3.2.2), $(ii)$ translate these into validity conditions for the admissible range of values of $\theta_{\mathrm{c}}$ within the hSIDM scheme (see Sect. 3.2.3), $(iii)$ demonstrate the resulting speed-up of the numerical algorithm based on hSIDM, and $(iv)$ provide a stringent validation test for our numerical implementation (see Sect. 5 below). Moreover, the test set-up can also be seen as a proxy how momentum and energy is transferred in a collision of two “patches” of the dark matter population within the two halos that collide in the initial stages of a galaxy cluster merger, provided the relative velocity is large compared to the velocity dispersion with each of the patches.
@@ 3.2.1 Deflection test set-up
The approximation given by Eq. (24) is the essential assumption underlying the effective treatment of small-angle scattering entering the fSIDM approach. This can be seen more directly when assuming in addition that $\tau_{\mathrm{T}}$ is small enough such that sufficiently many $\ell$ contribute in the summation in Eq. (25). In this case the summation over $\ell$ can be replaced by a continuous integration using the Euler formula (see e.g. Bethe, 1953), and Eq. (25) yields the Molière approximation (compare with Molière, 1948)
@@ 3.2.2 Validity of the fSIDM approach
Importantly, the analytically solvable deflection set-up allowed us to derive a general criterion for the validity of the effective forward-dominated approximation. In this subsection, we start by discussing the validity in the case of allowing for the entire range of possible scattering angles, i.e. the fSIDM approximation. Subsequently, we generalise the discussion to the hSIDM case, for which only angles below a critical angle are included in the effective small-angle approximation.
@@
where we expressed the number density in terms of the energy density using $\rho=m_{\chi}n$. For the effective small-angle approach to be useful, $t_{\rm min}$ should be much smaller than the timescale on which the distribution reaches its late-time isotropic limit, which is of order of the typical dynamical evolution timescale $t_{\rm dyn}\equiv 1/(\rho v\frac{\sigma_{\mathrm{T}}}{m_{\chi}})$, where $v$ is the beam velocity. Thus, we arrive at the condition for the effective small-angle treatment to be applicable and useful, given by
@@
We note that this is the condition that is relevant when using the effective fSIDM approach for all possible deflection angles, being in practice limited to models that exclusively feature forward-dominated scattering. For realistic models, this is typically not the case. As an example, Møller scattering with $r=1,10,100,10^{3},10^{4}$ yields $\sigma_{\mathrm{T}^{2}}/\sigma_{\mathrm{T}}=0.32,0.23,0.14,8.5\times 10^{-2},5.9\times 10^{-2}$ respectively. Analogously Rutherford scattering yields $1.2,0.82,0.51,0.33,0.24$. This motivates the use of the hybrid approach instead. We discuss its validity next.
@@ 3.2.3 Validity of the hSIDM approach
where $\tau_{X,<}(t,\theta_{\mathrm{c}})\equiv nvt\sigma_{X,<}(\theta_{\mathrm{c}})$ for $X=\mathrm{T}$ (transfer cross section) and $X=T^{2}$ (transfer-squared cross section, defined analogously to Eq. (29) but integrated only over the angle-interval $I_{<}$). As above, this implies a minimal timescale after which the hSIDM approach is valid, given by
@@
The conditions from above can be easily generalised to the case of identical particles with equal mass by using the appropriate definition of the cross sections $\sigma_{X,<}(\theta_{\mathrm{c}})$ for that case, (a) integrated over the interval corresponding $I_{<}=[0,\theta_{\mathrm{c}}]\cup[\uppi-\theta_{\mathrm{c}},\uppi]$ instead of $I_{<}=[0,\theta_{\mathrm{c}}]$, and (b) replacing the transfer ($X=\rm T$) by the modified transfer ($X=\rm{\widetilde{T}}$) cross section, see Sect. 2.2, and analogously for $X=\mathrm{T}^{2}$.
As an illustrative example, for Møller (Rutherford) scattering with $r=10^{3}$, one obtains $\sigma_{\mathrm{T}^{2},<}(\theta_{\mathrm{c}})/\sigma_{\mathrm{T},<}(\theta_{\mathrm{c}})=1.3\times 10^{-3}\,(2.6\times 10^{-3})$ for $\theta_{\mathrm{c}}=0.1$. We provide a detailed comparison of hSIDM simulations to a full, explicit treatment of the angle dependence for various models in Sect. 5, where we also cross-check the theoretically expected validity conditions derived here.
We expect that the condition on the cross section Eq. (36) is valid also beyond the deflection set-up. We note that it is independent of the details of the test problem and only involves the angle dependence of the cross section itself, and thus Eq. (36) depends only on the properties of the particle physics model underlying the angle dependence of the self-scattering cross section. Moreover, as discussed above, the deflection test can be viewed as a proxy for the dynamics of two colliding patches of dark matter in a galaxy cluster merger, such that the main physical mechanism of self-interactions on the dark matter distribution is broadly similar. This supports the applicability of Eq. (36) in cluster merger simulation set-ups. We also expect that Eq. (35) can be generalised beyond the test problem when replacing $t_{\text{dyn}}$ on the right-hand side by the appropriate dynamical timescale on which the system of interest evolves.